OpenStudy (anonymous):

The length of a rectangle is 4 centimeters less than twice its width. Find the dimensions if the area of the rectangle is 96 square centimeters.

OpenStudy (anonymous):

8 and 12

OpenStudy (anonymous):

I could've guessed that, but without showing work I'm afraid my teacher will give me no credit. I've tried setting it up algebraically but can't seem to get the correct answer. L= 2w-4cm w= w (2w-4)(w)=96 Don't know what I have wrong so if anyone could tell me it'd be appreciated.

OpenStudy (anonymous):

"length is 4 cm less than twice its width" L = 2W - 4 cm "area is 96 square centimeters" LW = 96 cm² (2W-4)W = 96 2W² - 4W = 96 2W² - 4W - 96 = 0 Use the quadratic formula to solve aW²+bW+c=0 where a=2 b=-4 c=-96 W = [-b ± √(b²-4ac)] / (2a) = [-(-4) ± √((-4)² - 4(2)(-96))] / 2(2) = [4 ± √(784)] / 4 = [4 ± √(2⁴)√(7²)] / 4 = [4 ± 4×7] / 4 = [4 ± 28] / 4 = 1 ± 7 W₊ = 1 + 7 = 8 W₋ = 1 - 7 = -6 invalid W = 8 cm L = 2W-4 = 12 cm

OpenStudy (anonymous):

Thanks!